3lout:

4x + 1 over 3y2 What is the value of the expression above when x = 2 and y = 3? You must show all work and calculations to receive full credit.

3lout:

btw this is math

3lout:

jit help me plsssssss

YRJ8498:

is that and exponent as the 2? example: does it look like this \[3y ^{2}\]

3lout:

yea

3lout:

i think

3lout:

can you answer this

SemiDefinite:

you can do it

3lout:

capp

iosangel:

sub x for 2 and y for 3 and solve like normal

ReturnofQuestion:

I agree.

carmelle:

I'm late on this, but... \[\frac{ 4x+1 }{ 3y ^{2} }\] All you have to do is substitute the x and y values with the numbers given and solve. x is 2 y is 3 \[\frac{ 4(2)+1 }{ 3(3) ^{2} }\] \[\frac{ 8+1 }{ 3(9) }=\frac{ 9 }{ 27 }=\frac{ 1 }{ 3 }\] Your final answer is 1/3. Written steps: 1. Substitute the x and y values. 2. Multiply 4 and 2. 3. Do the exponent, which is basically 3 × 3 4. Add 8 and 1 5. Multiply 3 and 9. 6. Reduce the fraction by dividing the numerator and denominator by 9.

3lout:

u were to late but thats ok

3lout:

and you were wrong

carmelle:

Oh, may I ask where I went wrong?

3lout:

the answer

3lout:

it was 11

3lout:

but don't worry bout it

carmelle:

Oh! That seems like a whole other equation. Next time, please post what the actual equation looks like because you had me thinking it was \[\frac{ 4x+1 }{ 3y ^{2} }\] when it was actually...\[4x + \frac{ 1 }{ 3 }y ^{2}\] if I'm not mistaken (?) I'm really sorry it was wrong though!

BLUEFACE123:

Putting the values in equation for x =2 and for y =3 4(2)+1/3×3×2 =8+1/3×6 8+6/3 10